# GraphOracle: A Foundation Model for Knowledge Graph Reasoning

### v1

@@ 4.2 Relation Representation Learning on Relation-Dependency Graph
Following Eq. (1). we apply a labeling initialization to distinguish the query relation node $r_{q}$ in $\mathcal{G}^{\mathcal{R}}$. Then employ multi-head attention relation-dependency message passing over the graph:
$\begin{array}[]{ll}&\bm{h}_{r_{v}|r_{q}}^{0}=\texttt{INDICATOR}_{r}(r_{v},r_{q})=\delta_{r_{v},r_{q}}\cdot\bm{1}^{d},\ \ \ r_{v}\in\mathcal{G}^{\mathcal{R}}\\[5.0pt] &\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\left(\frac{1}{H}\sum_{h=1}^{H}\left[\bm{W}_{1}^{\ell,h}\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\right]\right),\end{array}$ (5)
where $\delta_{r_{v},r_{q}}=1$ if $v=q$, and 00 otherwise. $H$ is the number of attention heads, and $\bm{W}^{\ell,h}_{1},\bm{W}^{\ell,h}_{2}\in\mathbb{R}^{d\times d}$ are head-specific parameter matrices. The relation-dependency attention weight $\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}$ captures the directional influence of relation $r_{u}$ on relation $r_{v}$, computed as:
$\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}=\frac{\exp(\bm{a}^{T}[\bm{W}^{h}_{\alpha}\bm{h}_{r_{u}}^{\ell-1}\|\bm{W}_{\alpha}^{h}\bm{h}_{r_{v}}^{\ell-1}])}{\sum_{r_{w}\in\mathcal{N}^{\text{past}}(r_{v})}\exp(\bm{a}^{T}[\bm{W}_{\alpha}^{h}\bm{h}_{r_{w}}^{\ell-1}\|\bm{W}_{\alpha}^{h}\bm{h}_{r_{v}}^{\ell-1}])},$ (6)
[…]
@@ 4.3 Entity Representation Learning on the Original KG
After obtaining the relation representations $\bm{R}_{q}$ for $r_{q}$ from RDG, we obtain query-dependent entity representations by conducting message passing over the original KG structures. This approach enables effective reasoning across both seen and unseen entities and relations.
@@
$\begin{array}[]{ll}\bm{h}^{\ell}_{e|q}=\delta\left(\bm{W}^{\ell}\cdot\sum_{(e_{s},r,e)\in\mathcal{F}_{\text{train}}}\alpha^{\ell}_{e_{s},r|q}\left(\bm{h}^{\ell-1}_{e_{s}|q}+\bm{h}^{L_{r}}_{r|r_{q}}\right)\right),\end{array}$ (7)
@@
$\alpha_{e_{s},r|q}^{\ell}=\sigma\Big{(}(\bm{w}_{\alpha}^{\ell})^{\top}\text{ReLU}\big{(}\bm{W}_{\alpha}^{\ell}\cdot(\bm{h}^{\ell-1}_{e_{s}|q}\|\bm{h}^{L_{r}}_{r|r_{q}}\|\bm{h}^{L_{r}}_{r_{q}|r_{q}})\big{)}\Big{)}.$ (8)
where $\bm{w}_{\alpha}^{\ell}\in\mathbb{R}^{d}$ and $\bm{W}_{\alpha}^{\ell}\in\mathbb{R}^{d\times 3d}$ are learnable parameters, and $\sigma$ is the sigmoid function [^1] Theoretical analysis of the GraphOracle model is given in Appendix H, and time complexity is given in Appendix B.1..
We iterate Eq. (7) for $L_{e}$ steps and use the final layer representation $\bm{h}_{e|q}^{L_{e}}$ for scoring each entity $e\in\mathcal{V}$. The critical idea here is replacing the learnable relation embeddings $\bm{r}$ with the contextualized relation embedding $\bm{h}^{L_{r}}_{r|r_{q}}$ from our relation-dependency relation graph, enabling fully inductive reasoning.
@@ 4.4 Training Details
All the learnable parameters such as $\bigl{\{}\bm{W}^{h}_{O},\;\bm{W}^{\ell,h}_{1},\;\bm{W}^{\ell,h}_{2},\;\bm{W}^{h}_{\alpha},\;\bm{a},\;\bm{W}^{\ell},\;\bm{W}^{\ell}_{\alpha},\;\bm{w}^{\ell}_{\alpha},\;\bm{w}^{L}\bigr{\}}$ are trained end-to-end by minimizing the loss function Eq. (2). GraphOracle adopts a sequential multi-dataset pre-train $\rightarrow$ fine-tune paradigm to acquire a general relation-dependency graph representation across KGs $\{\mathcal{G}_{1},\dots,\mathcal{G}_{K}\}$. For each graph $\mathcal{G}_{k}$, we optimize the regularized objective $\mathcal{L}^{(k)}=\mathcal{L}_{\text{task}}^{(k)}+\lambda_{k}\,\bigl{\lVert}\Theta\bigr{\rVert}_{2}^{\;2},$ where $\mathcal{L}_{\text{task}}^{(k)}$ denotes the task-specific loss on $\mathcal{G}_{k}$, $\Theta$ represents all learnable parameters, and $\lambda_{k}$ controls the strength of L2 regularization. Early stopping technique is used for each graph once validation MRR fails to improve for several epochs. This iterative pre-train process, together with our relation-dependency graph encoder, equips GraphOracle with strong cross-domain generalization. (Full datails are reported in Appendix D.) When adapting GraphOracle to new KGs, we firstly build the relation-dependency graph and then support two inference paradigms:
@@ 5.1 Experimental Setup
Datasets. We conduct comprehensive experiments on $31$ KGs, which we classify into three categories according to their properties (The detail of these datasets can be seen in Appendix C):
  • Transductive datasets with fixed entities and relations: WN18RR [8], FB15k-237 [36], NELL-995 [44], and YAGO3-10 [32].
  • Inductive datasets with (i) Fixed relations with new entities. Following [34], we use 12 subsets derived from WN18RR, FB15k237, and NELL-995 datasets. (ii) New relations and entities. Following INGRAM [19], we use 12 newly subsets derived from above datasets.
  • Cross domain datasets. (i) Biomedical Datasets: We use biomedical KG PrimeKG [3] to examine the cross-domain capabilities of GraphOracle. We finetune with 80% samples in raw PrimeKG, and validate with 10% samples. When testing on the remaining 10%, we specially focus on the predictions for triplets: (Protein, Interacts_with, BP/MF/CC), (Drug, Indication, Disease), (Drug, Target, Protein), (Protein, Associated_with, Disease), (Drug, Contraindication, Disease). (ii) Recommendation domain: We transform the Amazon-book [41] dataset into a pure KG reasoning dataset to adapt to the KGs Reasoning field by defining the interactions between users and items as a new relation in the KG. (iii) Geographic datasets (GeoKG) [15] . (The detail process of cross domain dataset can be found in Appendix E.)
Pretrain and Finetune. GraphOracle is pre-trained on three general KGs (NELL-995, FB15k-237, YAGO3-10) to capture diverse relational structures and reasoning patterns. We train GraphOracle for 150,000 steps with batch size of 32 using AdamW optimizer [24] on a single A6000 (48GB) GPU. For cross-domain adaptation, we employ a lightweight fine-tuning approach that updates only the final layer parameters while freezing the pre-trained representations. The finetune process only takes $1\sim 2$ epochs to achieve the best results. The pre-training process takes approximately 36 hours, while fine-tuning requires only $15\sim 60$ minutes depending on the target dataset. Detailed hyperparameters, architecture specifications, and training configurations are provided in Appendix D.
Baselines. We compare the proposed GraphOracle with (i) Transductive: ConvE [7], QuatE [47], RotatE [33], DuASE [20] and BioBRIDGE [43]; (ii) Entity inductive: MINERVA [6], DRUM [28], AnyBURL [26], RNNLogic [27], RLogic [4] GraphRulRL [25], CompGCN [38], NBFNet [58], RED-GNN [51], A*Net [56], Adaprop [52] and one-shot-subgraph [55]; (iii) Fully inductive: INGRAM [19] and ULTRA [11]. The results for the baseline methods were either directly obtained from the original publications or reproduced using the official source code provided by the authors. Due to page limitations, some other baselines can be found in INGRAM [19], BioBRIDGE [43] and KUCNet [22].
@@ 5.2 Overall Performance (RQ1)
(a) Nell-100
(b) WK-100
(c) FB-100
Figure 4: Learning Curve Comparison of GraphOracle (Fine-tuned vs. Trained from Scratch) against Supervised SOTA on Fully Inductive Datasets.
The main experimental results, presented in Fig. 3, illustrate the performance of GraphOracle following pre-training on three KGs and a brief fine-tuning phase of just two epochs across 31 distinct datasets (comprehensive results are available in Appendix F). A salient finding is that GraphOracle consistently outperforms the supervised SOTA across all evaluated baseline datasets. This robust performance underscores the overall efficacy of our methodology, culminating in an average MRR improvement of 35.48%. When disaggregated by reasoning type, this substantial gain is reflected in average improvements of 15.62% on transductive datasets, 21.99% on entity-inductive datasets, a significant 68.76% on fully inductive datasets, and 29.87% on cross-domain datasets.
To further investigate the benefits of our pre-training and fine-tuning strategy, particularly in challenging fully inductive scenarios, we analyzed the learning trajectories of GraphOracle. Fig. 4 depicts these learning curves on three classic fully-inductive KGs. The results highlight several key observations: Firstly, the pre-trained GraphOracle exhibits strong zero-shot performance (at epoch 0) and achieves remarkable gains with minimal fine-tuning (typically within 1-3 epochs). This is in stark contrast to the SOTA baseline, ULTRA-finetune, which tends to show degrading performance after its initial fine-tuning phase on these fully inductive tasks. Secondly, while GraphOracle trained from scratch eventually reaches competitive, SOTA-level performance, it consistently underperforms the fine-tuned version throughout the training process and requires significantly more epochs. This comparative analysis underscores the profound impact of our pre-training strategy; it enables GraphOracle to acquire a deeper and more efficient structural understanding and richer relational representations, which are crucial for rapid adaptation and superior generalization across novel relational structures.
5.3 Relation-Dependency Pattern Analysis (RQ2)
To investigate whether GraphOracle truly internalizes the compositionality of the relations, that is, the way complex relations are constructed systematically from simpler ones, we performed a series of perturbation analyzes on the learned RDG. First, we calculated relation attention weights using Eq. (6), averaged over the WN18RR and YAGO3-10 datasets, to assign an importance score to each relation pair, reflecting its contribution to downstream reasoning. Subsequently, during inference, we systematically disabled specific subsets of edges based on these attention weights: (i) the top-5 and top-10 most important (highest attention) compositional relation pairs; (ii) the bottom-5 and bottom-10 least important (lowest attention) pairs; and (iii) 5 and 10 randomly selected pairs.
5.4 Compatible with Additional Initial Information (RQ3)
[…]
To explore the potential of external information in enhancing KG reasoning, we introduced an improved entity initialization strategy. This involved incorporating modality-specific encoded features as initial entity vectors, moving beyond standard random initialization. The resulting model, denoted as GraphOracle+, leverages foundation model embeddings to create more semantically rich entity representations (details are provided in Appendix G). Experimental evaluations on the PrimeKG dataset show that GraphOracle+ achieves consistent performance gains across all metrics. Notably, MRR scores improved by 15% for protein-biological process prediction and 14% for protein-molecular function prediction. These results affirm that GraphOracle’s framework significantly benefits from integrating external information. In an era increasingly influenced by large language models, the capacity for flexible incorporation of diverse information sources is crucial for advancing generalization and adaptability, especially within specialized and complex domains such as biomedicine.
5.5 Ablation Study (RQ4)
@@
To rigorously evaluate the contribution of each architectural component within GraphOracle, we conducted an extensive series of ablation experiments. We first investigates the impact of removing the RDG and the effect of reducing the number of attention heads $H$ in Eq. (5) from eight to one. The results are detailed in Table 2. Clearly, eliminating the RDG or the multi-head attention mechanism causes a marked decline in all evaluation metrics, highlighting their indispensibility to GraphOracle’s performance.
In addition, we quantified how the breadth of pre-training data affects zero-shot performance. Specifically, we pre-trained on one to six heterogeneous datasets and evaluated the resulting checkpoints on unseen Table 2’s graphs. The averaged results, depicted in Fig. 6 (implementation details and complete scores are reported in Table 5 and Table 14), reveal that zero-shot performance saturates once four diverse datasets are included in the pre-training mixture. Incorporating additional datasets beyond this point yields no further significant gains. We conjecture that after this point the model has already encountered a sufficiently rich spectrum of relational patterns, and subsequent datasets may introduce largely redundant or potentially noisy signals.

### v2

@@ 4.2 Relation Representation Learning on RDG
Following Eq. (1). we apply a labeling initialization to distinguish the query relation node $r_{q}$ in $\mathcal{G}^{\mathcal{R}}$. Then employ multi-head attention relation‑dependency message passing over the graph:
$\bm{h}_{r_{v}\mid r_{q}}^{0}=\texttt{INDICATOR}_{r}(r_{v},r_{q})=\delta_{r_{v},r_{q}}\cdot\bm{1}^{d},\quad r_{v}\in\mathcal{G}^{\mathcal{R}}$ (4)
$\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\Big(\frac{1}{H}\sum_{h=1}^{H}\Big[\quad\;\bm{W}_{1}^{\ell,h}\!\!\!\!\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}$
$\quad\quad\quad\quad\quad+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\Big]\Big),$
where $\delta_{r_{v},r_{q}}=1$ if $v=q$, and 0 otherwise. $H$ is the number of attention heads, and $\bm{W}^{\ell,h}_{1},\bm{W}^{\ell,h}_{2}\in\mathbb{R}^{d\times d}$ are head-specific parameter matrices. The relation‑dependency attention weight $\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}$ captures the directional influence of relation $r_{u}$ on relation $r_{v}$, computed as:
$\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}=\frac{\exp\big(\bm{a}^{T}(\bm{W}^{h}_{\alpha}\bm{h}_{r_{u}}^{\ell-1}\|\bm{W}_{\alpha}^{h}\bm{h}_{r_{v}}^{\ell-1})\big)}{\sum_{r_{w}\in\mathcal{N}^{\text{past}}(r_{v})}\exp\big(\bm{a}^{T}(\bm{W}_{\alpha}^{h}\bm{h}_{r_{w}}^{\ell-1}\|\bm{W}_{\alpha}^{h}\bm{h}_{r_{v}}^{\ell-1})\big)},$
where $\bm{a}\in\mathbb{R}^{2d}$ is a learnable attention parameter vector, $\|$ denotes vector concatenation, and $\bm{W}^{h}_{\alpha}\in\mathbb{R}^{d\times d}$ are head-specific trainable projection matrix. The neighborhood function $\mathcal{N}^{\text{past}}(r_{v})$ enforces the relation‑dependency ordering of relations as defined in Eq. (3).
@@ 4.3 Entity Representation Learning on the Original KG
After obtaining the relation representations $\bm{R}_{q}$ from RDG conditioned on $r_{q}$, we obtain query-dependent entity representations by conducting message passing over the original KG structures. This approach enables effective reasoning across both seen and unseen entities and relations.
@@
$\begin{array}[]{ll}\!\!\!\!\bm{h}^{\ell}_{e|q}=\delta\!\left(\bm{W}^{\ell}\!\cdot\!\sum_{(e_{s},r,e)\in\mathcal{F}_{\text{train}}}\!\alpha^{\ell}_{e_{s},r|q}\!\big(\bm{h}^{\ell-1}_{e_{s}|q}+\bm{h}^{L_{r}}_{r|r_{q}}\big)\!\right)\!,\!\!\end{array}$ (5)
@@
$\alpha_{e_{s},r|q}^{\ell}=\sigma\Big((\bm{w}_{\alpha}^{\ell})^{\top}\text{ReLU}\big(\bm{W}_{\alpha}^{\ell}\cdot(\bm{h}^{\ell-1}_{e_{s}|q}\|\bm{h}^{L_{r}}_{r|r_{q}}\|\bm{h}^{L_{r}}_{r_{q}|r_{q}})\big)\Big).$
where $\bm{w}_{\alpha}^{\ell}\in\mathbb{R}^{d}$ and $\bm{W}_{\alpha}^{\ell}\in\mathbb{R}^{d\times 3d}$ are learnable parameters, $\sigma$ is the sigmoid function and $\cdot$ denotes the standard matrix-vector multiplication.
We iterate Eq. (5) for $L_{e}$ steps and use the final layer representation $\bm{h}_{e|q}^{L_{e}}$ for scoring each entity $e\in\mathcal{V}$. The critical idea here is replacing the learnable relation embeddings $\bm{r}$ with the contextualized relation embedding $\bm{h}^{L_{r}}_{r|r_{q}}$ from our RDG, enabling fully inductive reasoning (Time complexity of the GraphOracle model is given in Appendix C, and the Theoretical analysis on its expressiveness and generalization is given in Appendix I).
@@ 4.4 Training Details
All the learnable parameters such as $\bigl\{\bm{W}^{h}_{O},\;\bm{W}^{\ell,h}_{1},\;\bm{W}^{\ell,h}_{2},\;\bm{W}^{h}_{\alpha},\;\bm{a},\;\bm{W}^{\ell},\;\bm{W}^{\ell}_{\alpha},\;\bm{w}^{\ell}_{\alpha},\;\bm{w}^{L}\bigr\}$ are trained end-to-end by minimizing the loss function Eq. (2). GraphOracle adopts a sequential multi‑dataset pre‑train $\rightarrow$ fine‑tune paradigm to acquire a general relation-dependency graph representation across KGs $\{\mathcal{G}_{1},\dots,\mathcal{G}_{K}\}$. For each graph $\mathcal{G}_{k}$, we optimize the regularized objective $\mathcal{L}^{(k)}=\mathcal{L}_{\text{task}}^{(k)}+\lambda_{k}\,\bigl\lVert\Theta\bigr\rVert_{2}^{\;2},$ where $\mathcal{L}_{\text{task}}^{(k)}$ denotes the task‑specific loss on $\mathcal{G}_{k}$ (e.g., Eq. (2)), $\Theta$ represents all learnable parameters, and $\lambda_{k}$ controls the strength of L2 regularization. Early stopping technique is used for each graph once validation MRR fails to improve for several epochs. This iterative pre‑train process, together with our relation‑dependency graph encoder, equips GraphOracle with strong cross‑domain generalization. When adapting GraphOracle to new KGs, we firstly build the RDG and then support two inference paradigms:
@@
Model Transductive Entity Inductive Fully Inductive Cross-domain
Metric MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
Supervised SOTA 0.4185 0.4715 0.5771 0.4915 0.4730 0.6296 0.4593 0.2942 0.6246 0.2964 0.2149 0.4458
GraphOracle 0.4486 0.5550 0.6111 0.5449 0.5684 0.6722 0.5203 0.3688 0.7279 0.3759 0.2744 0.5485
Improvement 7.19% 17.70% 5.89% 10.86% 20.16% 6.77% 13.28% 25.36% 16.54% 26.82% 27.70% 23.03%
Table 2: Average performance comparison between GraphOracle and Supervised SOTA under four generalization settings.
@@ 5.1 Experimental Setup
Datasets
We conduct comprehensive experiments on $60$ KGs, which we classify into three categories according to their properties (Details are given in Appendix D.):
  • Transductive and Inductive datasets. To ensure fair comparison, we follow the same dataset settings as ULTRA (galkin2024ultra), TRIX (zhang2025trix), and KG-ICL (cui2024promptkg), including 16 transductive, 18 entity-inductive, and 23 fully-inductive datasets, totaling 57 in all.
  • Cross domain datasets. (i) Biomedical Datasets: We use biomedical KG PrimeKG (chandak2023building) to examine the cross-domain capabilities of GraphOracle. We finetune with 80% samples in raw PrimeKG, and validate with 10% samples. When testing on the remaining 10%, we specially focus on the predictions for triplets: (Protein, Interacts_with, BP/MF/CC), (Drug, Indication, Disease), (Drug, Target, Protein), (Protein, Associated_with, Disease), (Drug, Contraindication, Disease). (ii) Recommendation domain: We transform the Amazon-book (wang2019kgat) dataset into a pure KG reasoning dataset to adapt to the KGs Reasoning field by defining the interactions between users and items as a new relation in the KG. (iii) Geographic datasets (GeoKG) (geonames) . (Detail process are given in Appendix F.)
Pretrain and Finetune.
GraphOracle is pre-trained on three general KGs (NELL-995, CoDEx-Medium, FB15k-237) to capture diverse relational structures and reasoning patterns. It takes 150,000 training steps with batch size of 32 using AdamW optimizer (loshchilov2019decoupled) on a single A6000 (48GB) GPU. For cross-domain adaptation, we employ a lightweight fine-tuning approach that updates only the final layer parameters while freezing the pre-trained representations. The finetune process only takes $1\sim 2$ epochs to achieve the best results. The pre-training process takes approximately 36 hours, while fine-tuning requires only $15\sim 60$ minutes depending on the target dataset. Detailed hyperparameters, architecture specifications, and training configurations are provided in Appendix E.
Baselines
We compare the proposed GraphOracle with (i) Transductive: ConvE (ConvE), QuatE (QuatE), DuASE (li2024duase) and BioBRIDGE (wang2024biobridge); (ii) Entity inductive: MINERVA (MINERVA), DRUM (DRUM), AnyBURL (meilicke2020reinforced), RNNLogic (RNNLogic), RLogic (RLogic) GraphRulRL (mai2025graphrulrl), CompGCN (CompGCN), NBFNet (zhu2021neural), RED-GNN (RED-GNN), A*Net (Zhu2023AStarNet), Adaprop (Adaprop) and one-shot-subgraph (one-shot-subgraph); (iii) Fully inductive: INGRAM (lee2023ingram), ULTRA (galkin2024ultra), TRIX (zhang2025trix) and KG-ICL (cui2024promptkg). The results for the baseline methods were either directly obtained from the original publications or reproduced using the official source code provided by the authors. Due to page limitations, some other baselines can be found in INGRAM (lee2023ingram), BioBRIDGE (wang2024biobridge) and KUCNet (liu2024kucnet).
@@ 5.2 Overall Performance (RQ1)
Figure 5: Comparison on PrimeKG: Evaluating GraphOracle Enhanced by External Entity information (GraphOracle+).
@@
The main experimental results, presented in Fig. 2, illustrate the performance of GraphOracle following pre-training on three KGs and a brief fine-tuning phase of just two epochs across 60 distinct datasets (comprehensive results are available in Appendix G). A salient finding is that GraphOracle consistently outperforms the supervised SOTA across all evaluated baseline datasets and metrics, as detailed in Table 2. This robust performance underscores the overall efficacy of our methodology, with particularly notable improvements observed in the more challenging scenarios. The results demonstrate substantial gains across all reasoning types, with the most pronounced improvements occurring in fully inductive and cross-domain settings, where the model must generalize to entirely unseen entities or domains—scenarios that represent the most stringent tests of a model’s reasoning capabilities.
5.3 Relation-Dependency Pattern Analysis (RQ2)
To investigate whether GraphOracle truly internalizes the compositionality of the relations, that is, the way complex relations are constructed systematically from simpler ones, we performed a series of perturbation analyzes on the learned RDG. First, we calculated relation attention weights using Eq. (4.2), averaged over the WN18RR and YAGO3-10 datasets, to assign an importance score to each relation pair, reflecting its contribution to downstream reasoning. Subsequently, during inference, we systematically disabled specific subsets of edges based on these attention weights: (i) the top-5 and top-10 most important (highest attention) compositional relation pairs; (ii) the bottom-5 and bottom-10 least important (lowest attention) pairs; and (iii) 5 and 10 randomly selected pairs.
5.4 Compatible with Additional Initial Information (RQ3)
To explore the potential of external information in enhancing KG reasoning, we introduced an improved entity initialization strategy. This involved incorporating modality-specific encoded features as initial entity vectors, moving beyond standard random initialization. The resulting model, denoted as GraphOracle+, leverages foundation model embeddings to create more semantically rich entity representations (details are provided in Appendix H). As demostrated in Fig 5 (details are given in Tabel 17), experimental evaluations on the PrimeKG dataset show that GraphOracle+ achieves consistent performance gains across all metrics. Notably, MRR scores improved by 15% for protein-biological process prediction and 14% for protein-molecular function prediction. These results affirm that GraphOracle’s framework significantly benefits from integrating external information. In an era increasingly influenced by large language models, the capacity for flexible incorporation of diverse information sources is crucial for advancing generalization and adaptability, especially within specialized and complex domains such as biomedicine.
5.5 Ablation Study (RQ4)
To rigorously evaluate the contribution of each architectural component within GraphOracle, we conducted an extensive series of ablation experiments. We first investigate the impact of removing the RDG and the effect of reducing the number of attention heads $H$ in Eq. (4) from eight to one. The results are detailed in Table 3. Clearly, eliminating the RDG or the multi‑head attention mechanism causes a marked decline in all evaluation metrics, highlighting their indispensibility to GraphOracle’s performance.
In addition, we quantified how the breadth of pre‑training data affects zero‑shot performance. Specifically, we pre‑trained on one to six heterogeneous datasets and evaluated the resulting checkpoints on unseen Table 3’s graphs. The averaged results, depicted in Fig. 4 (implementation details are reported in Appendix E), reveal that zero-shot performance saturates once three diverse datasets are included in the pre-training mixture. Incorporating additional datasets beyond this point yields no further significant gains. We conjecture that after this point the model has already encountered a sufficiently rich spectrum of relational patterns, and subsequent datasets may introduce largely redundant or potentially noisy signals.